PLEASE SHOW ALL WORK!

A given RL circuit has a characteristic time constant of 99.0 ns and a resistance of 40.0
kΩ.

a) What is the inductance of the circuit?

b) What is the current in the circuit after two time constants have elapsed if it’s
connected to a high-voltage source of 160 kV?
Answer #1

time Constant = 2 = L/R Z R=L 19 3 L - 99 x 10 x 40 x 10 L- -3 3-9 6x 10 H. the - 22/ 그 wil E e 160 40 (t=22) 0 I = u e I=0-5